How much pure acid should be mixed with 2 gallons of a 40% acid solution in order to get an 80%acid solution

Answer by stanbon(75887)

How much pure acid should be mixed with 2 gallons of a 40% acid solution in order to get an 80% acid solution
-------------------
Equation:
acid + acid = acid
0.40*2 + 1*x = 0.80(2+x)
----
Multiply thru by 100 to get:
40*2 + 100x = 80*2 + 80x
20x = 40*2
x = 4 gallons (amt. of pure acid needed)
===============
Cheers,
Stan H.
===============